The Ksp of Al(OH)3 is 2 × 10–32. At what pH will a 0.4 M Al3+ solution begin to show precipitation of Al(OH)3?

6.1

1.0

3.6

3.1

10.4
Answer #1

Al(OH)3 <------------> Al+3 + 3OH-

Ksp = [Al+3] [OH-]3

[Al+3] = S moles /L

[OH-] = 3S moles /L

2.0 x 10-32 = [0.4] [3S]3

2.0 x 10-32 = 0.4 x 27 S3

2.0 x 10-32 = 10.8 S3

S3 = 1.85 x 10-33

S = 1.23 x 10-11

[OH-] = 3S = 3 x 1.23 x 10-11 = 3.69 x 10-11M

pOH = - log [OH-]

pOH = - log [3.69 x 10-11]

pOH = 10.4

pH = 14 - 10.4

pH = 3.6

answer = 3.6
Answer #2
2×10^-32=[0.4][27x^3] 2×10^-32/10.8=x^3 X=1.227×10^-11 pH=-log of x And:10.9
source: Zumdah zumdahl 9 th edition
answered by: anonymous